40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 9638     Entry time: Fri Feb 14 02:33:09 2014
Author: manasa 
Type: Update 
Category: General 
Subject: My IFO time summary 

MC tuning

Although the morning MC tuning looked stable, Koji pointed out that the MC_REFL_OFFSET was changed from its nominal value.

The offset was reset and this caused drift in the MC_TRANS_SUM.

To fix this:

- disabled the WFS servo

- aligned MC using MC1 and MC3

- centered beam on the MC_REFL

- reset WFS offsets

- locked MC

MC looks happy now.

__________________________________________________________________________________________________________________

ALS locking

ALS is in a very different state from a couple of days ago when we could successfully lock the arms and scan.

The green alignment to the arms had drifted.

PSL green alignment on the PSL table was off. The PSL green was not even on the steering mirror. Did anyone work around the PSL table in the last couple of days?

After aligning and finding the beat note, I found the ALS servo very noisy. The error signal had 10 times more rms noise than what was achieved earlier this week and there were some new 60Hz peaks as well.

Overall, we could not do any PRMI+ALS arms today

 

 

 

 

ELOG V3.1.3-